Suppose a 12-foot ladder is leaning against a wall so that its base is 2 feet from the edge of the wall. As you climb to the very top of the ladder, you find you can reach 7 feet above the top of the ladder. To what height are you reaching(in feet, to one decimal place)?
Answer by rfer(16322)

a^2=c^2-b^2
a^2=12^2-2^2
a^2=144-4
a^2=140
a=11.8
11.8+7=18.8 ft reach